An object of mass m is dropped from height h above
Answer #1

Mass of the object m = m
Mass of the planet M = M
Radius of the planet R = R
Height h = h
Total energy at height h is E = Potential energy at height h
                                          = -GMm/(R+h)
Total energy at the surface of the planet E ' = [-GMm/R] + ( 1/ 2) mv 2
From law of conservation of energy , E = E '

-GMm/(R+h) =  [-GMm/R] + ( 1/ 2) mv 2
( 1/ 2) mv 2 = [-GMm/(R+h) ] +[GMm/R]

             v 2 = [-2GM/(R+h) ] +[2GM/R]

                  = [-2GMR +2GM(R+h)]/[(R+h)R]

                 = 2GMh / [(R+h) R ]

             v = √ {2GMh / [(R+h) R ]}